OpenStudy (anonymous):

Please help me! Complete the table below giving the amount P that must be invested at interest rate 7 % compounded weekly to obtain a balance of A = $40000 in t years. t=1;10;20;30;40;50 i got the correct values for t=1;10; and 20 but for some reason using the exact same formula i get the wrong answers for the rest. The formula i used was A=Pe^rt

OpenStudy (ranga):

No. You use A=Pe^rt when the interest is compounded continuously. Here it is compounded weekly and so you have to use a different formula.

OpenStudy (ranga):

Here is the compound interest formula: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\]A = Amount at maturity P = Principal Amount r = Annual interest rate in decimal n = compounding period (compounded how many times a year) t = years invested

OpenStudy (anonymous):

ohhh.. then it is just a coincidence that it worked for the ones i got right?

OpenStudy (ranga):

yes!

OpenStudy (ranga):

Here it is compounded weekly and therefore n = 52 because there are 52 weeks in a year. A = 40,000 ; r = 0.07 ; t=1;10;20;30;40;50 Calculate the corresponding Ps

OpenStudy (anonymous):

okay. Thank you so much!

OpenStudy (ranga):

you are welcome.